OpenStudy (anonymous):

How do i find the two real irrational solutions of this equation? G(x)=x^2+14x+18

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hellp please im so lost

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

do you know quadratic formula ?

hartnn (hartnn):

Compare your quadratic function with \(G(x)= ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two solutions of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

i'm still super confused...sorry

hartnn (hartnn):

could you first find a,b,c only? \(x^2+14x+18\) \(ax^2+bx+c\)

OpenStudy (anonymous):

so a=1 b=14 and c=18?

hartnn (hartnn):

yes! b^2-4ac = ... ?

OpenStudy (anonymous):

196-72?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes!

OpenStudy (anonymous):

what next? :)

hartnn (hartnn):

just plug the values in the formula given

OpenStudy (anonymous):

so plug a=1 b=14 and c=18 into the fraction thing you showed me? sorry i'm botherin u :/

hartnn (hartnn):

yes, no prob and you already found b^2-4ac in it

OpenStudy (anonymous):

so -4|sqrt sign 124 over 2?

hartnn (hartnn):

\((-14 \pm \sqrt{124})/2 \)

hartnn (hartnn):

and you can simplify that

OpenStudy (anonymous):

........im sorry i'm not sure i can :/

hartnn (hartnn):

124 = 4*31

hartnn (hartnn):

(-7 \(\pm\sqrt{31}\))

OpenStudy (anonymous):

how did you find the 7?

hartnn (hartnn):

-14/2 = -7

OpenStudy (anonymous):

right.sorry so the answer has to be 2 solutions but the sign b4  means it can either b positive or negative so thats two solutions right?

hartnn (hartnn):

yes, \( -7-\sqrt{31} \\ -7+\sqrt{31}\)

OpenStudy (anonymous):

thanks so much!

hartnn (hartnn):

welcome ^_^